# HydroGEM: A Self-Supervised Zero Shot Hybrid TCN-Transformer Foundation Model for Continental Scale Streamflow Quality Control

### v2

@@ 4.4 Evaluation Protocols
For synthetic evaluation we use a held out set of 799 USGS sites that are completely unseen during training. Clean sequences from these sites are corrupted with injected anomalies that cover a broad range of types, durations, spatial coverage, and severities. We report aggregate detection metrics over all anomalies and also stratify performance by anomaly type and basic regime attributes such as duration and coverage. Detailed breakdowns are provided in the appendix, while the main text focuses on overall detection performance and the relative ranking of HydroGEM against baselines.
@@
Canadian zero shot evaluation.
For cross national evaluation we apply HydroGEM without retraining to 100 Canadian stations from Environment and Climate Change Canada. Weak labels are derived from human corrections in the operational record. These labels reflect editorial decisions and local correction practice, so we report F1 and related detection metrics over a range of labeling thresholds defined by the relative magnitude of human edits, with full curves in the appendix. This provides a sensitivity view of how performance changes as the notion of meaningful correction becomes stricter. Given the documented differences between USGS and Environment and Climate Change Canada correction strategies in ice affected periods [41, 44, 18], we interpret reconstruction errors here as supportive evidence rather than a primary outcome and focus our conclusions on detection performance.
@@ 5.1.1 Overall Detection Performance
Table 6 and Figure 3 present detection performance for HydroGEM and eleven baseline methods on the synthetic test set. HydroGEM achieves F1 = 0.792 (precision: 0.755, recall: 0.832), substantially outperforming all baselines. The strongest baseline, Isolation Forest, attains F1 = 0.392, representing a 101.9% relative improvement by HydroGEM. All pairwise improvements are statistically significant ($p<0.001$, Wilcoxon signed-rank test).
Table 6: Detection performance on the synthetic test set. All baseline thresholds were derived from site-specific data statistics following QARTOD methodology, ensuring no parameter tuning on test data.
@@
The baseline hierarchy reveals distinct performance tiers. Isolation Forest (F1 = 0.392) achieves the strongest baseline performance by leveraging multivariate structure in the discharge-stage feature space. Traditional statistical methods (Z-Score, IQR, Moving Average) achieve F1 << 0.15, as they operate on univariate signals and cannot capture stage-discharge coupling violations. Domain-specific approaches like Q-H Consistency (F1 = 0.118) and Rating Curve residuals (F1 = 0.056) underperform despite their hydrological grounding—these methods detect only anomalies that violate instantaneous rating relationships, missing temporally-extended distortions that preserve local Q-H ratios.
@@ 5.1.2 Per-Anomaly-Type Analysis
Performance varies systematically with anomaly characteristics. HydroGEM achieves strongest detection for backwater effects (F1 = 0.77), ice backwater (F1 = 0.76), and drift anomalies (F1 = 0.73)—conditions that produce sustained Q-H decoupling patterns. Moderate performance occurs for sensor-level artifacts including bias steps (F1 = 0.61), flatlines (F1 = 0.63), and spikes (F1 = 0.61), which manifest as localized discontinuities rather than extended temporal patterns. The most challenging anomaly type is rating shift (F1 = 0.51), where abrupt but physically-plausible changes in the stage-discharge relationship can resemble legitimate rating curve updates.
@@ 5.1.3 Reconstruction Quality
Beyond detection, HydroGEM provides corrected time series through its reconstruction head. On the synthetic test set, the model achieves 68.7% mean error reduction, computed as:
$\text{Error Reduction}=1-\frac{\text{MAE}(\hat{x},x_{\text{clean}})}{\text{MAE}(x_{\text{corrupt}},x_{\text{clean}})}$ (2)
where $\hat{x}$ denotes the HydroGEM reconstruction, $x_{\text{clean}}$ the ground-truth clean signal, and $x_{\text{corrupt}}$ the corrupted input.
@@ 5.1.4 Representative Detection Examples
Figure 6 presents a challenging case containing multiple gate operation events producing rapid, physically-valid discharge fluctuations. HydroGEM achieves perfect recall (1.000) but lower precision (0.622), resulting in F1 = 0.767 with slightly negative error reduction (-3.2%). The model flags legitimate gate-induced transients as suspicious—these rapid Q-H excursions resemble sensor artifacts in the learned representation. This case illustrates the inherent difficulty of distinguishing anthropogenic flow modifications from sensor malfunctions without explicit metadata about control structures.
Figure 6: Challenging case with gate operation events. HydroGEM achieves F1 = 0.767 with perfect recall but reduced precision, as legitimate gate-induced transients produce Q-H patterns similar to sensor artifacts. This motivates treating model outputs as suggestions requiring expert review.
@@ 5.2 Zero-Shot Transfer to Canadian Stations
The ultimate test of a foundation model is generalization beyond its training distribution. We evaluated HydroGEM on 100 stations from Environment and Climate Change Canada’s (ECCC) hydrometric network—data the model never encountered during training. This assessment tests whether representations learned from USGS stations capture universal principles of streamflow monitoring that transfer across national boundaries, instrumentation practices, and climatic regimes.
@@ 5.2.1 Baseline Comparison
Table 7 presents detection performance on Canadian stations. We applied the same eleven baseline methods using identical QARTOD-derived threshold protocols described in Section 5.1.1. HydroGEM achieves F1 = 0.586 at the 1% correction threshold, substantially outperforming all baselines with improvements statistically significant at $p<0.001$ (Wilcoxon signed-rank test).
Table 7: Detection performance on Canadian zero-shot evaluation (F1 at 1% correction threshold).
@@
HydroGEM Foundation Model 0.586 0.544 0.635
@@
The baseline ranking differs notably between synthetic and Canadian evaluations. Isolation Forest performs strongly on synthetic data (F1 = 0.392) but less so on Canadian data (F1 = 0.263), while Persistence shows the opposite pattern (synthetic: F1 = 0.001; Canadian: F1 = 0.418). This divergence reflects differences in anomaly composition: synthetic anomalies include diverse corruption patterns where multivariate outlier detection excels, whereas Canadian operational corrections predominantly address sensor stalls and ice effects that Persistence can partially identify through its stuck-sensor detection. HydroGEM’s consistent strong performance across both evaluation settings shows robustness to these distributional differences.
5.2.2 Representative Detection Example
@@
The ice-affected period (March 16–24, blue shading) exhibits reconstruction divergence where ECCC operators applied aggressive discharge reductions while stage remained elevated—a characteristic ice-backwater correction strategy. HydroGEM’s proposed reconstructions reflect discharge-stage relationships learned from USGS training data, which differ systematically from Canadian practices in ice-control periods. This divergence does not indicate detection failure; rather, it shows that the model learned USGS correction principles rather than memorizing site-specific patterns. For Canadian deployment, we recommend treating reconstructions as suggestions requiring operator review, particularly during ice-affected periods.
@@ 5.2.3 Detection Behavior Characterization
We compared HydroGEM’s flag rate against the human correction rate across seasons (Figure 8A). Both exhibit consistent seasonal structure: winter shows the highest activity (HydroGEM: 54%, human: 68%), followed by spring (51%, 60%), fall (43%, 56%), and summer (39%, 42%). The winter peak reflects well-documented challenges in cold-region hydrometry: ice-affected flow produces backwater conditions violating standard rating curves, anchor ice and frazil ice cause sensor interference, and instrument freezing creates data gaps [34, 54, 6]. That HydroGEM—trained on predominantly ice-free USGS stations—identifies these signatures in Canadian data suggests the model learned generalizable representations of sensor malfunction and rating curve violation.
@@
A concern with learned detectors is whether they simply flag large deviations while missing subtle corrections. Figure 8B stratifies detection recall by correction magnitude. HydroGEM maintains consistent recall (0.48–0.56) across corrections ranging from minor adjustments (1–5%) to major revisions (50–100%), indicating the model learned anomaly patterns—temporal dynamics, channel inconsistencies, physical constraint violations—rather than relying on deviation magnitude as a proxy.
@@
5.3 Summary of Results
Across both evaluation settings, HydroGEM shows strong anomaly detection performance with substantial margins over established baselines. On synthetic data with exact ground truth, the model achieves F1 = 0.792 with 68.7% reconstruction error reduction, outperforming the best baseline by 36.3%. On Canadian stations in zero-shot transfer, HydroGEM achieves F1 = 0.586, exceeding the best baseline by 40.2%. The model generalizes across all 18 expert-validated anomaly types, aligns with operational correction patterns across seasons, and maintains sensitivity to both subtle and severe anomalies.
The architectural separation between training anomalies (simplified patterns in normalized space) and test anomalies (physically-grounded injection with multiple equation variants) confirms that these results reflect learned hydrometric principles rather than pattern memorization. Zero-shot transfer to Canadian stations—with different instrumentation, correction practices, and ice-dominated hydrology—provides additional evidence of generalizable representations. Complete per-window results and visualizations are available at https://github.com/[REPO]/hydrogem-results.
@@ 5.4.2 Architectural Component Analysis
Figure 10: Architectural ablation results. (a) Detection performance by anomaly duration shows TCN-only degradation beyond its 61-hour receptive field, while Transformer-only fails on short events. (b) Training curves demonstrate 2x slower convergence without gated skip connections.
@@
Conversely, the Transformer-only variant shows opposite behavior: poor performance on short transients (F1 = 0.45 for 1-6 hour events) but improving with duration. Self-attention struggles to isolate brief spikes within 576-hour windows, effectively smoothing high-frequency patterns that constitute many operational anomalies.
The gated skip connection implements crucial curriculum learning. Without it, convergence requires approximately 20 epochs versus 10 with the skip path, and final reconstruction error remains 35% higher. The learnable gate parameter evolves from near-zero initially ($\sigma(\alpha)\approx 0.007$) to moderate values ($\approx 0.35$) by training completion, progressively incorporating learned representations while maintaining gradient flow.
@@ 5.4.3 Training Strategy Validation
Two-stage (ours) 0.684 ± 0.021 96.8% 12 epochs
End-to-end synthetic 0.521 ± 0.034 89.3% 8 epochs
No pretraining 0.398 ± 0.048 91.2% 15+ epochs
Pretrain only (no finetune) 0.216 ± 0.019 98.1% N/A
@@
The pretrained backbone provides stable initialization with learned representations of recession curves, rating relationships, and seasonal patterns. This foundation enables rapid fine-tuning convergence and superior final performance while maintaining 96.8% clean data preservation—critical for operational trust.
@@ 5.4.5 Synthesis of Ablation Findings
These ablations demonstrate that HydroGEM’s performance emerges from synergistic interaction between components rather than any single innovation:
  • Hierarchical normalization allows learning across six orders of magnitude while preserving scale-aware detection capabilities
  • Hybrid architecture captures complementary temporal scales, with TCN handling short transients and Transformer modeling extended dependencies
  • Two-stage training prevents overfitting to synthetic patterns while establishing robust hydrological priors
  • Multi-component loss ensures physically plausible reconstructions meeting operational requirements
Performance degrades 20-42% when removing major components, confirming the current architecture represents a minimal configuration for continental-scale deployment. Each design choice addresses specific failure modes encountered when scaling anomaly detection beyond individual sites to thousands of heterogeneous monitoring stations.
@@ 6.1 Interpretation of Key Findings
The successful development and validation of HydroGEM shows that foundation model principles can be effectively adapted to operational hydrological quality control, addressing a critical bottleneck in environmental monitoring infrastructure. The results reveal several important insights about both the model’s capabilities and the nature of streamflow data quality challenges.
The two-stage training approach proves remarkably effective at learning generalizable hydrometric principles without extensive labeled anomaly datasets. By pretraining on 6.03 million clean sequences, HydroGEM develops robust representations of normal hydrological behavior that serve as an inductive bias for anomaly detection. Recession curves, rating relationships, and seasonal patterns emerge naturally from the self-supervised objective. The subsequent fine-tuning with simplified synthetic anomalies forces the model to learn fundamental consistency principles rather than memorizing specific corruption signatures. The four-axis separation between training and test distributions provides strong evidence that the 36.3% performance gain over baselines reflects genuine abstraction rather than pattern matching. This finding suggests that the historical challenge of limited labeled anomaly data in operational hydrology can be circumvented through careful architectural design and training strategies.
The hierarchical normalization scheme successfully allows learning across six orders of magnitude in discharge, from ephemeral desert washes to major rivers, within a single model. This addresses a fundamental challenge in continental-scale hydrology where traditional approaches either require site-specific calibration or suffer from large-river dominance. The combination of logarithmic stabilization, site-specific standardization, and scale embeddings preserves both local patterns and absolute magnitude information. The model can thus distinguish a 10 ft3/s anomaly in a small creek from a 10,000 ft3/s anomaly in a major river, despite their identical appearance in normalized space. This capability is essential for continental-scale deployment where manual calibration of thousands of sites would be prohibitively expensive.
Zero-shot transfer to Canadian stations shows that learned representations capture universal principles of streamflow monitoring rather than USGS-specific artifacts. The alignment between HydroGEM’s detection patterns and operational correction rates across seasons is particularly striking. The winter peak in both model flagging and human corrections, corresponding to documented ice effects [54, 6], emerges despite training predominantly on ice-free stations. This suggests the model learned to identify fundamental signatures of sensor malfunction and rating curve violation that manifest similarly across different hydrological contexts. The ability to generalize across national boundaries, instrumentation protocols, and climatic regimes validates the foundation model approach for operational hydrology.

### v3

@@ 4.4 Evaluation Protocols
For synthetic evaluation we use the held-out set of 799 USGS sites described in Section 2.3.1 that are completely unseen during training. Clean sequences from these sites are corrupted with injected anomalies that cover a broad range of types, durations, spatial coverage, and severities. We report aggregate detection metrics over all anomalies and also stratify performance by anomaly type and basic regime attributes such as duration and coverage. Detailed breakdowns are provided in the appendix, while the main text focuses on overall detection performance and the relative ranking of HydroGEM against baselines.
@@
Canadian zero-shot evaluation.
For cross-national evaluation we apply HydroGEM without retraining to 100 Canadian stations from Environment and Climate Change Canada. Weak labels are derived from human corrections in the operational record as described in Section 2.3.2. These labels reflect editorial decisions and local correction practice, so we report tolerant F1 (with $\pm$24-hour buffer) as the primary metric, alongside pointwise F1 and segment-level recall, over a range of weak-label thresholds defined by the relative magnitude of human edits, with full curves in the appendix. This provides a sensitivity view of how performance changes as the notion of meaningful correction becomes stricter. Given the documented differences between USGS and Environment and Climate Change Canada correction strategies in ice-affected periods [54, 58, 20], we interpret reconstruction errors here as supportive evidence rather than a primary outcome and focus our conclusions on detection performance.
@@ 5.1.1 Overall Detection Performance
[…]
Table 7: Detection performance on the synthetic test set. All baseline hyperparameters and thresholds are fixed a priori (Appendix J) and are not tuned against synthetic test labels.
@@
The baseline hierarchy reveals distinct performance tiers. Isolation Forest (F1 = 0.392) achieves the strongest baseline performance by leveraging multivariate structure in the discharge-stage feature space. Traditional statistical methods (Z-Score, IQR, Moving Average) achieve F1 << 0.15, as they operate on univariate signals and cannot capture stage-discharge coupling violations. Domain-specific approaches like Q–H Consistency (F1 = 0.118) and Rating Curve residuals (F1 = 0.056) underperform despite their hydrological grounding—these methods detect only anomalies that violate instantaneous rating relationships, missing temporally extended distortions that preserve local Q–H ratios.
@@ 5.1.2 Per-Anomaly-Type Analysis
Performance varies systematically with anomaly characteristics. HydroGEM achieves strongest detection for backwater effects (F1 = 0.77), ice backwater (F1 = 0.76), and drift anomalies (F1 = 0.73)—conditions that produce sustained Q–H decoupling patterns. Moderate performance occurs for sensor-level artifacts including bias steps (F1 = 0.61), flatlines (F1 = 0.63), and spikes (F1 = 0.61), which manifest as localized discontinuities rather than extended temporal patterns. The most challenging anomaly type is rating shift (F1 = 0.51), where abrupt but physically plausible changes in the stage-discharge relationship can resemble legitimate rating curve updates.
@@ 5.1.3 Reconstruction Quality
Beyond detection, HydroGEM provides suggested corrections through its reconstruction output. On the synthetic test set, the model achieves 68.7% mean error reduction computed using the metric defined in Section 4.2.
@@ 5.1.4 Detection Examples
Figure 7 presents a challenging case containing multiple gate operation events producing rapid, physically valid discharge fluctuations. HydroGEM achieves recall of 1.0 but lower precision (0.622), resulting in F1 = 0.767 with slightly negative error reduction (-3.2%). The model flags legitimate gate-induced transients as suspicious—these rapid Q–H excursions resemble sensor artifacts in the learned representation. This case illustrates the inherent difficulty of distinguishing anthropogenic flow modifications from sensor malfunctions without explicit metadata about control structures.
Figure 7: Challenging case with gate operation events. HydroGEM achieves F1 = 0.767 with high recall but reduced precision, as legitimate gate-induced transients produce Q–H patterns similar to sensor artifacts. This motivates treating model outputs as suggestions requiring expert review.
@@ 5.2 Zero-Shot Transfer to Canadian Stations
We evaluated HydroGEM on 100 stations from Environment and Climate Change Canada’s (ECCC) hydrometric network, data the model never encountered during training. This assessment tests whether representations learned from USGS stations transfer across national boundaries, instrumentation practices, and climatic regimes.
@@ 5.2.1 Baseline Comparison
Table 8 presents pointwise detection performance on Canadian stations. We applied the same eleven baseline methods using identical threshold protocols described in Section 5.1.1. HydroGEM achieves pointwise F1 = 0.582, substantially outperforming all baselines (paired Wilcoxon signed-rank test over 100 sites, $p<0.001$).
Table 8: Pointwise detection performance on Canadian zero-shot evaluation (1% correction threshold).
@@
HydroGEM Foundation Model 0.582 0.650 0.567
@@
The baseline ranking differs notably between synthetic and Canadian evaluations. Isolation Forest performs strongly on synthetic data (F1 = 0.392) but less so on Canadian data (F1 = 0.263), while Persistence shows the opposite pattern (synthetic: F1 = 0.001; Canadian: F1 = 0.418). This divergence reflects differences in anomaly composition: synthetic anomalies include diverse corruption patterns where multivariate outlier detection excels, whereas Canadian operational corrections predominantly address sensor stalls and ice effects that Persistence can partially identify. HydroGEM’s consistent strong performance across both evaluation settings demonstrates robustness to these distributional differences.
5.2.2 Detection Example
@@
The ice-affected period (March 16–24, blue shading) exhibits reconstruction divergence where ECCC operators applied aggressive discharge reductions while stage remained elevated, a characteristic ice-backwater correction strategy. HydroGEM’s proposed reconstructions reflect discharge-stage relationships learned from USGS training data, which differ systematically from Canadian practices in ice-control periods. This divergence does not indicate detection failure; rather, it demonstrates that the model learned USGS correction principles rather than memorizing site-specific patterns. For Canadian deployment, reconstructions should be treated as suggestions requiring operator review, particularly during ice-affected periods.
5.2.3 Evaluation Under Weak Labels
The ECCC ground-truth labels are derived from operational corrections applied by Water Survey of Canada technicians. These corrections reflect practical data quality decisions rather than precise anomaly boundaries: a correction recorded as starting at midnight may correspond to an anomaly whose true onset occurred hours earlier. Standard pointwise F1 penalizes detections that correctly identify anomaly events but with minor temporal offset, a known limitation for time series anomaly detection on weakly labeled data [33].
To address this, we adopt evaluation metrics from the time series anomaly detection literature that account for temporal tolerance and interval-level detection. Tolerant F1 credits a prediction if it falls within $\pm$24 hours of a labeled correction timestep, accommodating the daily granularity typical of correction records [70]. Segment-level recall treats contiguous anomaly intervals as discrete events rather than independent timestamps and counts an event as detected if any predicted anomaly overlaps the interval or falls within the tolerance window [27].
Table 9 presents HydroGEM’s performance across these complementary metrics.
Table 9: Zero-shot HydroGEM performance on 100 ECCC stations (1% correction threshold).
Metric Precision Recall F1/Score
Pointwise F1 0.650 0.567 0.582
Tolerant F1 ($\pm$24h) 0.683 0.765 0.700
Segment-level recall 0.901
Figure 9: Illustration of tolerant evaluation on Canadian station 02CF011. Pointwise F1 (0.521) penalizes boundary misalignment despite correct event detection; tolerant F1 with $\pm$24-hour buffer (0.995) credits temporally proximate predictions.
We adopt Tolerant F1 = 0.70 as the primary metric for this evaluation based on three considerations. First, tolerant evaluation aligns with operational practice: technicians review flagged periods holistically rather than verifying individual timestamps, so detections within 24 hours of recorded corrections remain operationally valuable. Second, the $\pm$24-hour buffer accommodates inherent timing imprecision in correction records without being excessively permissive. Third, precision remains stable across all buffer sizes tested (0.654 at $\pm$1h to 0.683 at $\pm$24h), indicating that the tolerance does not artificially inflate performance but simply avoids penalizing temporally proximate detections.
At the segment level, HydroGEM detects 90.1% of anomaly events with at least partial overlap (segment recall = 0.901), demonstrating effective identification of anomaly intervals rather than just individual points. Extended evaluation including tolerance sensitivity analysis, weighted F1, and range score is provided in Appendix I.
@@ 5.2.4 Detection Behavior Characterization
We compared HydroGEM’s flag rate against the human correction rate across seasons (Figure 10A). Both exhibit consistent seasonal structure: winter shows the highest activity (HydroGEM: 54%, human: 68%), followed by spring (51%, 60%), fall (43%, 56%), and summer (39%, 42%). The winter peak reflects well-documented challenges in cold-region hydrometry: ice-affected flow produces backwater conditions violating standard rating curves, anchor ice and frazil ice cause sensor interference, and instrument freezing creates data gaps [87, 78, 9]. That HydroGEM, trained on predominantly ice-free USGS stations, identifies these signatures in Canadian data suggests the model learned generalizable representations of sensor malfunction and rating curve violation rather than site-specific patterns.
@@
A concern with learned detectors is whether they simply flag large deviations while missing subtle corrections. Figure 10B stratifies detection recall by correction magnitude. HydroGEM maintains consistent recall (0.48–0.56) across corrections ranging from minor adjustments (1–5%) to major revisions (50–100%), indicating the model learned anomaly patterns based on temporal dynamics, channel inconsistencies, and physical constraint violations rather than relying on deviation magnitude as a proxy.
@@ 5.3.2 Architectural Component Analysis
[…]
@@
Conversely, the Transformer-only variant shows opposite behavior: poor performance on short transients (F1 = 0.45 for 1–6 hour events) but improving with duration. Self-attention struggles to isolate brief spikes within 576-hour windows, effectively smoothing high-frequency patterns that constitute many operational anomalies.
The gated skip connection facilitates stable early training. Without it, convergence requires approximately 20 epochs versus 10 with the skip path, and final reconstruction error remains 35% higher. The gate parameter $\alpha$ is learned during training, starting near zero and increasing as the model converges, progressively incorporating learned representations while maintaining gradient flow.
@@ 5.3.3 Training Strategy Validation
Two-stage (Ours) 0.684 $\pm$ 0.021 96.8% 12 epochs
End-to-end synthetic 0.521 $\pm$ 0.034 89.3% 8 epochs
No pretraining 0.398 $\pm$ 0.048 91.2% 15+ epochs
Pretrain only (no Stage 2) 0.216 $\pm$ 0.019 98.1% N/A
@@
The pretrained backbone provides stable initialization with learned representations of recession curves, rating relationships, and seasonal patterns. This foundation enables rapid Stage 2 convergence and superior final performance while maintaining 96.8% clean data preservation—critical for operational trust.
@@ 5.3.5 Synthesis of Ablation Findings
These ablations demonstrate that HydroGEM’s performance emerges from synergistic interaction between components rather than any single innovation. Hierarchical normalization allows learning across six orders of magnitude while preserving scale-aware detection capabilities. The hybrid architecture captures complementary temporal scales, with TCN handling short transients and Transformer modeling extended dependencies. Two-stage training prevents overfitting to synthetic patterns while establishing robust hydrological priors. The multi-component loss ensures physically plausible reconstructions meeting operational requirements.
Performance degrades 20–42% when removing major components, confirming the current architecture represents a minimal configuration for continental-scale deployment. Each design choice addresses specific failure modes encountered when scaling anomaly detection beyond individual sites to thousands of heterogeneous monitoring stations.
@@ 6.1 Interpretation of Key Findings
HydroGEM demonstrates that foundation model principles can be effectively adapted to operational hydrological quality control, addressing a critical bottleneck in environmental monitoring infrastructure. The results reveal several important insights about both the model’s capabilities and the nature of streamflow data quality challenges.
[…]
The hierarchical normalization scheme successfully enables learning across six orders of magnitude in discharge, from ephemeral desert washes to major rivers, within a single model. This addresses a fundamental challenge in continental-scale hydrology where traditional approaches either require site-specific calibration or suffer from large-river dominance. The combination of logarithmic stabilization, site-specific standardization, and scale embeddings preserves both local patterns and absolute magnitude information. The model can thus distinguish anomalies in small creeks from those in major rivers, despite their similar appearance in normalized space. This capability is essential for continental-scale deployment where manual calibration of thousands of sites would be prohibitively expensive.
Zero-shot transfer to Canadian stations demonstrates that learned representations capture general principles of streamflow monitoring rather than USGS-specific artifacts. The alignment between HydroGEM’s detection patterns and operational correction rates across seasons is particularly notable. The winter peak in both model flagging and human corrections, corresponding to documented ice effects [78, 9], emerges despite training predominantly on ice-free stations. This suggests the model learned to identify fundamental signatures of sensor malfunction and rating curve violation that manifest similarly across different hydrological contexts. The ability to generalize across national boundaries, instrumentation protocols, and climatic regimes supports the foundation model approach for operational hydrology.